Let X and Y be random losses with joint density function and 0 otherwise. An insurance policy is written to reimburse X +Y. Calculate the probability that the reimbursement is less than 1.
Answer #1

here probability=P(X+Y<1)=\int_{0}^{1}\int_{0}^{1-x} f(x,y) dy dx =\int_{0}^{1}\int_{0}^{1-x}e-(x+y) dy dx =\int_{0}^{1}-e-x*e-y |1-x0 dx

=\int_{0}^{1}e-x(1-e-1+x) dx=\int_{0}^{1}(e-x-e-1) dx =(-e-x-xe-1)|10 =0.2642